Anna is now y years old, and Marco is 3 years younger. In terms of y, how old was Marco 5 years ago?

Answer by ankor@dixie-net.com(22740)

Anna is now y years old,
a = y
and Marco is 3 years younger.
m = y - 3
In terms of y, how old was Marco 5 years ago?
f(y) = m
:
f(y) = (y-3) - 5
f(y) = y - 8